if a man has a mass of 74 kilograms runs 800 meters in 150 seconds, what is his average momentum?
1
Expert's answer
2019-12-17T09:53:17-0500

The speed of the man


"v=\frac{d}{t}=\frac{800 \:\rm m}{150\:\rm s}=5.3\:\rm m/s"

The average momentum of the man


"p=mv=74\:\rm kg\times 5.3\: m/s=394.7\: kg\cdot m/s"